Pick out the solution from the values given in the bracket next to each equation. Show that the other values do not satisfy the equation.
r – 4 = 0                    (4, -4, 8, 0)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

r–4 = 0
r = 4 is a solution for this equation because for r = 4
r–4 =4–4=0
Equation satisfied
r = -4 is not a solution for this equation because for r = - 4,
r–4 =-4–4
= -8 and not 0
r = 8 is not a solution for this equation because for r = 8
r–4 =8–4
= 4 and not 0
r = 0 is not a solution for this equation because for r = 0,
r–4 =0–4
= - 4 and not 0

Thanks